### v1

@@ 1 Introduction
Таблица 1: Comparison of meanings and pronunciations using diacritics. The table illustrates differences in phonetic transcriptions and their associated meanings, highlighting how diacritics affect pronunciation and meaning.
@@
Таблица 2: illustrates how diacritics in Arabic can change based on the context of subsequent words in a sentence, dramatically altering meaning and grammatical function
@@
Таблица 3: Punctuation can cause diacritization ambiguities in Arabic by altering sentence structure and grammatical roles, as seen in shifts between case endings and changes in word function (e.g., verb vs. adjective).
@@ 4 Benchmark
Рис. 1: Benchmark Category Statistics
@@
By conducting the diacritization process in-house and incorporating expert review, we ensure that the dataset remains entirely novel to existing models, as they have not been exposed to the diacritized form of our data. This directly addresses the concern raised in the CATT Alasmary et al. [2024] regarding the potential familiarity of language models with benchmark datasets. Our approach effectively prevents data leakage and results in a genuinely unseen evaluation set, thereby enabling a more reliable and accurate assessment of models’ true diacritization capabilities. SadeedDiac-25 is publicly available [^1] https://huggingface.co/datasets/Misraj/Sadeed_Tashkeela to support further research and development in Arabic diacritization and to facilitate fair and reproducible evaluation of diacritization models.
@@ 5 Model
Рис. 2: Sadded pipeline, it shows how we get sadeed from our model kuwain-1.5
@@
Рис. 3: (a) is The System prompts we use to fine-tune our models. (b) An example from our dataset after applying the template
@@ 6.1 Overlap Analysis Between Fadel and Abbad Datasets
Таблица 4: Comparison of datasets showing the total number of samples, overlap of identical samples (with percentage), and overlap of samples with similarity greater than 0.5 (with percentage).
@@
Таблица 5: Comparison of datasets showing the total number of samples, overlap of identical samples (with percentage), and overlap of samples with similarity greater than 0.5 (with percentage).
@@ 7.1 Evaluation on Fadel Benchmark
Таблица 6: Performance comparison of Arabic diacritization models on the Fadel dataset in terms of Diacritic Error Rate (DER) and Word Error Rate (WER)
@@ 7.2 Evaluation on WikiNews Benchmark
Таблица 7: Performance comparison of Arabic diacritization models on the WikiNews dataset in terms of Diacritic Error Rate (DER) and Word Error Rate (WER)
@@ 7.3 Evaluation on the SadeedDiac-25 Benchmark
Таблица 8: Performance comparison of Arabic diacritization models on the SadeedDiac-25 dataset in terms of Diacritic Error Rate (DER) and Word Error Rate (WER)
@@ 8.4 Limitations
Список литературы
  • Abbad & Xiong [2020] Hamza Abbad and Shengwu Xiong. Multi-components system for automatic arabic diacritization. In Joemon M. Jose, Emine Yilmaz, João Magalhães, Pablo Castells, Nicola Ferro, Mário J. Silva, and Flávio Martins (eds.), Advances in Information Retrieval, pp.  341–355, Cham, 2020. Springer International Publishing. ISBN 978-3-030-45439-5.
  • Abdelali et al. [2016] Ahmed Abdelali, Kareem Darwish, Nadir Durrani, and Hamdy Mubarak. Farasa: A fast and furious segmenter for Arabic. In John DeNero, Mark Finlayson, and Sravana Reddy (eds.), Proceedings of the 2016 Conference of the North American Chapter of the Association for Computational Linguistics: Demonstrations, pp.  11–16, San Diego, California, June 2016. Association for Computational Linguistics. 10.18653/v1/N16-3003. URL https://aclanthology.org/N16-3003.
  • Al-Thubaity et al. [2020] Abdulmohsen Al-Thubaity, Atheer Alkhalifa, Abdulrahman Almuhareb, and Waleed Alsanie. Arabic diacritization using bidirectional long short-term memory neural networks with conditional random fields. IEEE Access, 8:154984–154996, 2020.
  • Alasmary et al. [2024] Faris Alasmary, Orjuwan Zaafarani, and Ahmad Ghannam. Catt: Character-based arabic tashkeel transformer, 2024. URL https://arxiv.org/abs/2407.03236.
  • AlKhamissi et al. [2020] Badr AlKhamissi, Muhammad N ElNokrashy, and Mohamed Gabr. Deep diacritization: Efficient hierarchical recurrence for improved arabic diacritization. arXiv preprint arXiv:2011.00538, 2020.
  • Almanea [2021] Manar M. Almanea. Automatic methods and neural networks in arabic texts diacritization: A comprehensive survey. IEEE Access, 9:145012–145032, 2021. 10.1109/ACCESS.2021.3122977.
  • Alnefaie & Azmi [2017] Rehab Alnefaie and Aqil M. Azmi. Automatic minimal diacritization of arabic texts. Procedia Computer Science, 117:169–174, 2017. ISSN 1877-0509. https://doi.org/10.1016/j.procs.2017.10.106. URL https://www.sciencedirect.com/science/article/pii/S1877050917321634. Arabic Computational Linguistics.
  • Aryabumi et al. [2024] Viraat Aryabumi, John Dang, Dwarak Talupuru, Saurabh Dash, David Cairuz, Hangyu Lin, Bharat Venkitesh, Madeline Smith, Jon Ander Campos, Yi Chern Tan, et al. Aya 23: Open weight releases to further multilingual progress. arXiv preprint arXiv:2405.15032, 2024.
  • Bahar et al. [2023] Parnia Bahar, Mattia Di Gangi, Nick Rossenbach, and Mohammad Zeineldeen. Take the hint: Improving arabic diacritization with partially-diacritized text, 2023. URL https://arxiv.org/abs/2306.03557.
  • Bari et al. [2024] M Saiful Bari, Yazeed Alnumay, Norah A Alzahrani, Nouf M Alotaibi, Hisham A Alyahya, Sultan AlRashed, Faisal A Mirza, Shaykhah Z Alsubaie, Hassan A Alahmed, Ghadah Alabduljabbar, et al. Allam: Large language models for arabic and english. arXiv preprint arXiv:2407.15390, 2024.
  • Darwish et al. [2017a] Kareem Darwish, Hamdy Mubarak, and Ahmed Abdelali. Arabic diacritization: Stats, rules, and hacks. In Nizar Habash, Mona Diab, Kareem Darwish, Wassim El-Hajj, Hend Al-Khalifa, Houda Bouamor, Nadi Tomeh, Mahmoud El-Haj, and Wajdi Zaghouani (eds.), Proceedings of the Third Arabic Natural Language Processing Workshop, pp.  9–17, Valencia, Spain, April 2017a. Association for Computational Linguistics. 10.18653/v1/W17-1302. URL https://aclanthology.org/W17-1302.
  • Darwish et al. [2017b] Kareem Darwish, Hamdy Mubarak, and Ahmed Abdelali. Arabic diacritization: Stats, rules, and hacks. In Nizar Habash, Mona Diab, Kareem Darwish, Wassim El-Hajj, Hend Al-Khalifa, Houda Bouamor, Nadi Tomeh, Mahmoud El-Haj, and Wajdi Zaghouani (eds.), Proceedings of the Third Arabic Natural Language Processing Workshop, pp.  9–17, Valencia, Spain, April 2017b. Association for Computational Linguistics. 10.18653/v1/W17-1302. URL https://aclanthology.org/W17-1302.
  • Darwish et al. [2020] Kareem Darwish, Ahmed Abdelali, Hamdy Mubarak, and Mohamed Eldesouki. Arabic diacritic recovery using a feature-rich bilstm model, 2020. URL https://arxiv.org/abs/2002.01207.
  • Diab et al. [2007] Mona Diab, Mahmoud Ghoneim, and Nizar Habash. Arabic diacritization in the context of statistical machine translation. In Proceedings of Machine Translation Summit XI: Papers, 2007.
  • Fadel et al. [2019a] Ali Fadel, Ibraheem Tuffaha, Bara’ Al-Jawarneh, and Mahmoud Al-Ayyoub. Arabic text diacritization using deep neural networks, 2019a. URL https://arxiv.org/abs/1905.01965.
  • Fadel et al. [2019b] Ali Fadel, Ibraheem Tuffaha, Bara’ Al-Jawarneh, and Mahmoud Al-Ayyoub. Neural arabic text diacritization: State of the art results and a novel approach for machine translation. In Proceedings of the 6th Workshop on Asian Translation, pp.  215–225. Association for Computational Linguistics, 2019b. 10.18653/v1/d19-5229. URL http://dx.doi.org/10.18653/v1/D19-5229.
  • Habash et al. [2016] Nizar Habash, Anas Shahrour, and Muhamed Al Khalil. Exploiting arabic diacritization for high quality automatic annotation. In Proceedings of the Tenth International Conference on Language Resources and Evaluation (LREC’16), pp.  4298–4304, 2016.
  • Hennara et al. [2025] Khalil Hennara, Sara Chrouf, Mohamed Motaism Hamed, Zeina Aldallal, Omar Hadid, and Safwan AlModhayan. Kuwain 1.5b: An arabic slm via language injection, 2025. URL https://arxiv.org/abs/2504.15120.
  • Kharsa et al. [2024] Ruba Kharsa, Ashraf Elnagar, and Sane Yagi. Bert-based arabic diacritization: A state-of-the-art approach for improving text accuracy and pronunciation. Expert Systems with Applications, 248:123416, 2024. ISSN 0957-4174. https://doi.org/10.1016/j.eswa.2024.123416. URL https://www.sciencedirect.com/science/article/pii/S0957417424002811.
  • Likic [2008] Vladimir Likic. The needleman-wunsch algorithm for sequence alignment. Lecture given at the 7th Melbourne Bioinformatics Course, Bi021 Molecular Science and Biotechnology Institute, University of Melbourne, pp.  1–46, 2008.
  • Maamouri et al. [2008] Mohamed Maamouri, Seth Kulick, and Ann Bies. Diacritic annotation in the Arabic treebank and its impact on parser evaluation. In Nicoletta Calzolari, Khalid Choukri, Bente Maegaard, Joseph Mariani, Jan Odijk, Stelios Piperidis, and Daniel Tapias (eds.), Proceedings of the Sixth International Conference on Language Resources and Evaluation (LREC’08), Marrakech, Morocco, May 2008. European Language Resources Association (ELRA). URL http://www.lrec-conf.org/proceedings/lrec2008/pdf/706_paper.pdf.
  • Madhfar & Qamar [2021] Mokthar Ali Hasan Madhfar and Ali Mustafa Qamar. Effective deep learning models for automatic diacritization of arabic text. IEEE Access, 9:273–288, 2021. 10.1109/ACCESS.2020.3041676.
  • Navid-AI [2025] Navid-AI. Yehia 7b preview. https://huggingface.co/Navid-AI/Yehia-7B-preview, 2025.
  • Nelken & Shieber [2005] Rani Nelken and Stuart M. Shieber. Arabic diacritization using weighted finite-state transducers. In Kareem Darwish, Mona Diab, and Nizar Habash (eds.), Proceedings of the ACL Workshop on Computational Approaches to Semitic Languages, pp.  79–86, Ann Arbor, Michigan, June 2005. Association for Computational Linguistics. URL https://aclanthology.org/W05-0711.
  • Pasha et al. [2014] Arfath Pasha, Mohamed Al-Badrashiny, Mona Diab, Ahmed El Kholy, Ramy Eskander, Nizar Habash, Manoj Pooleery, Owen Rambow, and Ryan Roth. MADAMIRA: A fast, comprehensive tool for morphological analysis and disambiguation of Arabic. In Nicoletta Calzolari, Khalid Choukri, Thierry Declerck, Hrafn Loftsson, Bente Maegaard, Joseph Mariani, Asuncion Moreno, Jan Odijk, and Stelios Piperidis (eds.), Proceedings of the Ninth International Conference on Language Resources and Evaluation (LREC’14), pp.  1094–1101, Reykjavik, Iceland, May 2014. European Language Resources Association (ELRA). URL http://www.lrec-conf.org/proceedings/lrec2014/pdf/593_Paper.pdf.
  • Sabtan [2021] Yasser Sabtan. Arabic part-of-speech tagging using a combined rule-based and data-driven approach. Digital Scholarship in the Humanities, 36(3):719–735, 2021.
  • Sengupta et al. [2023] Neha Sengupta, Sunil Kumar Sahu, Bokang Jia, Satheesh Katipomu, Haonan Li, Fajri Koto, William Marshall, Gurpreet Gosal, Cynthia Liu, Zhiming Chen, Osama Mohammed Afzal, Samta Kamboj, Onkar Pandit, Rahul Pal, Lalit Pradhan, Zain Muhammad Mujahid, Massa Baali, Xudong Han, Sondos Mahmoud Bsharat, Alham Fikri Aji, Zhiqiang Shen, Zhengzhong Liu, Natalia Vassilieva, Joel Hestness, Andy Hock, Andrew Feldman, Jonathan Lee, Andrew Jackson, Hector Xuguang Ren, Preslav Nakov, Timothy Baldwin, and Eric Xing. Jais and jais-chat: Arabic-centric foundation and instruction-tuned open generative large language models, 2023. URL https://arxiv.org/abs/2308.16149.
  • Skiredj & Berrada [2024] Abderrahman Skiredj and Ismail Berrada. Arabic text diacritization in the age of transfer learning: Token classification is all you need, 2024. URL https://arxiv.org/abs/2401.04848.
  • Team et al. [2024] Gemma Team, Morgane Riviere, Shreya Pathak, Pier Giuseppe Sessa, Cassidy Hardin, Surya Bhupatiraju, Léonard Hussenot, Thomas Mesnard, Bobak Shahriari, Alexandre Ramé, et al. Gemma 2: Improving open language models at a practical size. arXiv preprint arXiv:2408.00118, 2024.
  • Team [2024] Silma Team. Silma. 2024. URL https://www.silma.ai.
  • Zakraoui et al. [2021] Jezia Zakraoui, Moutaz Saleh, Somaya Al-Maadeed, and Jihad Mohamed Alja’am. Arabic machine translation: A survey with challenges and future directions. IEEE Access, 9:161445–161468, 2021.
  • Zerrouki & Balla [2017] Taha Zerrouki and Amar Balla. Tashkeela: Novel corpus of arabic vocalized texts, data for auto-diacritization systems. Data in Brief, 11, 02 2017. 10.1016/j.dib.2017.01.011.
@@
Таблица 9: Training Hyperparameters for Sadeed Model
@@
Приложение B Model Hallucinations Examples
@@
Таблица 10: Examples of hallucination correction in Arabic text. In Sentence 1, the grammatical case ending is corrected to align with the change in word type caused by the hallucinated word. In Sentences 2 and 3, the hallucinated words were foreign terms.
Приложение C Examples from Sadeed
@@
Таблица 11: Output samples from Sadeed
Приложение D CATT Benchmark Examples
@@
Таблица 12: In the table provided, the red color indicates the presence of a diacritical error in the word, while the green color denotes the corrected diacritical marking. The orange color represents cases where multiple valid forms are allowed depending on the context or according to different linguistic schools.

### v2

@@ 1 Introduction
Table 1: Comparison of meanings and pronunciations using diacritics. The table illustrates differences in phonetic transcriptions and their associated meanings, highlighting how diacritics affect pronunciation and meaning.
@@
Table 2: illustrates how diacritics in Arabic can change based on the context of subsequent words in a sentence, dramatically altering meaning and grammatical function
@@
Table 3: Punctuation can cause diacritization ambiguities in Arabic by altering sentence structure and grammatical roles, as seen in shifts between case endings and changes in word function (e.g., verb vs. adjective).
@@ 4 Benchmark
Figure 1: Benchmark Category Statistics
@@
By conducting the diacritization process in-house and incorporating expert review, we ensure that the dataset remains entirely novel to existing models, as they have not been exposed to the diacritized form of our data. This directly addresses the concern raised in the CATT Alasmary et al. [2024] regarding the potential familiarity of language models with benchmark datasets. Our approach effectively prevents data leakage and results in a genuinely unseen evaluation set, thereby enabling a more reliable and accurate assessment of models’ true diacritization capabilities. SadeedDiac-25 is publicly available [^1] https://huggingface.co/datasets/Misraj/SadeedDiac-25 to support further research and development in Arabic diacritization and to facilitate fair and reproducible evaluation of diacritization models.
@@ 5 Model
Figure 2: Sadded pipeline, it shows how we get sadeed from our model kuwain-1.5
@@
Figure 3: (a) is The System prompts we use to fine-tune our models. (b) An example from our dataset after applying the template
@@ 6.1 Overlap Analysis Between Fadel and Abbad Datasets
Table 4: Comparison of datasets showing the total number of samples, overlap of identical samples (with percentage), and overlap of samples with similarity greater than 0.5 (with percentage).
@@
Table 5: Comparison of datasets showing the total number of samples, overlap of identical samples (with percentage), and overlap of samples with similarity greater than 0.5 (with percentage).
@@ 7.1 Evaluation on Fadel Benchmark
Table 6: Performance comparison of Arabic diacritization models on the Fadel dataset in terms of Diacritic Error Rate (DER) and Word Error Rate (WER)
@@ 7.2 Evaluation on WikiNews Benchmark
Table 7: Performance comparison of Arabic diacritization models on the WikiNews dataset in terms of Diacritic Error Rate (DER) and Word Error Rate (WER)
@@ 7.3 Evaluation on the SadeedDiac-25 Benchmark
Table 8: Performance comparison of Arabic diacritization models on the SadeedDiac-25 dataset in terms of Diacritic Error Rate (DER) and Word Error Rate (WER)
@@ Appendix A Training Details
Table 9: Training Hyperparameters for Sadeed Model
@@
Appendix B Model Hallucinations Examples
@@
Table 10: Examples of hallucination correction in Arabic text. In Sentence 1, the grammatical case ending is corrected to align with the change in word type caused by the hallucinated word. In Sentences 2 and 3, the hallucinated words were foreign terms.
Appendix C Examples from Sadeed
@@
Table 11: Output samples from Sadeed
Appendix D CATT Benchmark Examples
@@
Table 12: In the table provided, the red color indicates the presence of a diacritical error in the word, while the green color denotes the corrected diacritical marking. The orange color represents cases where multiple valid forms are allowed depending on the context or according to different linguistic schools.